# An Improved Algorithm for Adversarial Linear Contextual Bandits via Reduction

### v1

@@ Appendix B Lemmas for Linear Programming
Lemma 3 (Corollaries 14.1f and  14.1g of Schrijver, (1986)).
@@
Lemma 4 (Corollary 14.1a of Schrijver, (1986)).
@@ Appendix C Proof of Lemma 1
We will first reduce the result to a statement about uniform convergence for linear multiclass classifiers with $K$ classes and an unusual loss function. To this end, note first that we can assume without loss of generality that $\mathcal{A}=\{V_{1},\ldots,V_{K}\}$, where the $V_{1},\ldots,V_{K}$ are random vectors in $\mathbb{R}^{d}$. (If $\mathcal{A}$ has fewer than $K$ elements, then consider it as a multiset and add repetitions of one of its elements.) We can assume all randomness in $\mathcal{A}$ is determined by an underlying random variable $Z$ and that $V_{y}=-g(Z,y)$ for each ‘class’ $y\in[K]$, where $g$ is a class-sensitive feature map in the sense of Shalev-Shwartz and Ben-David, (2014, Section 17.2). Defining the linear multiclass classifier
@@
and applying part 1 of Theorem 26.5 of Shalev-Shwartz and Ben-David, (2014) separately to $f$ and f-f to control both parts in the maximum separately using a union bound; then noting that $f$ and f-f have the same Rademacher complexity.)
@@
by Massart’s lemma (Shalev-Shwartz and Ben-David,, 2014). As discussed by Shalev-Shwartz and Ben-David, (2014, Chapter 29), one possible generalization of the Vapnik-Chervonenkis dimension to multiclass classification is Natarajan’s dimension $\operatorname*{NatDim}(\mathcal{C})$. Natarajan’s lemma (Natarajan,, 1989, p. 93), (Shalev-Shwartz and Ben-David,, 2014, Lemma 29.4) shows that
@@
and, for linear multiclass classifiers, it is also known (Shalev-Shwartz and Ben-David,, 2014, Theorem 29.7) that
@@ D.1 Robust Linear Bandits
Lemma 5 (Lemma 14 of Zimmert and Lattimore, (2022)).
@@
It is known that the continuous exponential weight algorithm is equivalent to FTRL with entropic barrier as the regularizer together with a particular sampling scheme (Bubeck and Eldan,, 2015; Zimmert and Lattimore,, 2022). We summarize the equivalence in the following lemma, the details of which can be found in Zimmert and Lattimore, (2022).
Lemma 6 (Facts from Bubeck and Eldan, (2015), Zimmert and Lattimore, (2022)).
@@
The analysis for the FTRL term follows that in Ito et al., (2020). Specifically, directly following their Lemma 5, Lemma 6, and the analysis below Lemma 6, we have the following: as long as $\eta\|\widehat{\theta}_{t}-b_{t}\|_{\Sigma_{t}}\leq 1$, we have

### v2

@@ Appendix B Lemmas for Linear Programming
Lemma 3 (Corollaries 14.1f and  14.1g of ilp_theory).
@@
Lemma 4 (Corollary 14.1a of ilp_theory).
@@
Lemma 5 (Theorem 3.1.2 of weibel2007minkowski).
Let $P_{1},\ldots,P_{N}$ be polytopes in $\mathbb{R}^{d}$. Let $F$ be a face of the Minkowski sum $\sum_{i=1}^{N}P_{i}$. Then there are faces $F_{1},F_{2},\ldots,F_{N}$ of $P_{1},P_{2},\ldots,P_{N}$ respectively such that $F=\sum_{i=1}^{N}F_{i}$. What’s more, this decomposition is unique.
Lemma 6 (Corollary 3.1.3 of weibel2007minkowski).
Let $P=\sum_{i=1}^{N}P_{i}$ be a Minkowsky sum of polytopes in $\mathbb{R}^{d}$, let $F$ be a nonempty face of $P$, and let $F_{1},\ldots,F_{N}$ be its decomposition. Then $\mathcal{N}(P,F)=\mathcal{N}(P_{1},F_{1})\cap\cdots\cap\mathcal{N}(P_{N},F_{N})$.
@@ Appendix C Proof of Lemma 1
We will first reduce the result to a statement about uniform convergence for linear multiclass classifiers with $K$ classes and an unusual loss function. To this end, note first that we can assume without loss of generality that $\mathcal{A}=\{V_{1},\ldots,V_{K}\}$, where the $V_{1},\ldots,V_{K}$ are random vectors in $\mathbb{R}^{d}$. (If $\mathcal{A}$ has fewer than $K$ elements, then consider it as a multiset and add repetitions of one of its elements.) We can assume all randomness in $\mathcal{A}$ is determined by an underlying random variable $Z$ and that $V_{y}=-g(Z,y)$ for each ‘class’ $y\in[K]$, where $g$ is a class-sensitive feature map in the sense of ShalevShwartzBenDavid2014. Defining the linear multiclass classifier
@@
and applying part 1 of Theorem 26.5 of ShalevShwartzBenDavid2014 separately to $f$ and f-f to control both parts in the maximum separately using a union bound; then noting that $f$ and f-f have the same Rademacher complexity.)
@@
by Massart’s lemma (ShalevShwartzBenDavid2014). As discussed by ShalevShwartzBenDavid2014, one possible generalization of the Vapnik-Chervonenkis dimension to multiclass classification is Natarajan’s dimension $\operatorname*{NatDim}(\mathcal{C})$. Natarajan’s lemma (Natarajan1989, p. 93), (ShalevShwartzBenDavid2014, Lemma 29.4) shows that
@@
and, for linear multiclass classifiers, it is also known (ShalevShwartzBenDavid2014, Theorem 29.7) that
@@ D.1 Robust Linear Bandits
Lemma 7 (Lemma 14 of zimmert2022return).
@@
It is known that the continuous exponential weight algorithm is equivalent to FTRL with entropic barrier as the regularizer together with a particular sampling scheme (bubeck2015entropic; zimmert2022return). We summarize the equivalence in the following lemma, the details of which can be found in zimmert2022return.
Lemma 8 (Facts from bubeck2015entropic, zimmert2022return).
@@
The analysis for the FTRL term follows that in ito2020tight. Specifically, directly following their Lemma 5, Lemma 6, and the analysis below Lemma 6, we have the following: as long as $\eta\|\widehat{\theta}_{t}-b_{t}\|_{\Sigma_{t}}\leq 1$, we have